OpenStudy (anonymous):

What are the coordinates of the inflection point on the graph of y=(x+1)arctanx a. (-1,0) b. (0,0) c. (0,1) d. (1, pi/4) e. (1, pi/2) work plzz :)

OpenStudy (amistre64):

2nd derivative is?

OpenStudy (anonymous):

i couldn't get that far :(

OpenStudy (anonymous):

i hav the first deriv

OpenStudy (amistre64):

we can go from there then, whats the 1st D ?

OpenStudy (anonymous):

f'(x)=1(1/1+x^2)

OpenStudy (amistre64):

hmmm, lets see if I can verify that: \[y=(x+1)tan^{-1}(x)\] \[y'=(x+1)'tan^{-1}(x)+(x+1)tan'^{-1}(x)\] \[y'=tan^{-1}(x)+\frac{x+1}{1+x^2}\] \[y'=tan^{-1}(x)+(x+1)(1+x^2)^{-1}\] you agree?

OpenStudy (anonymous):

ohhhhhhhhhhhh quotient rule!!

OpenStudy (amistre64):

\[y'=tan^{-1}(x)+(x+1)(1+x^2)^{-1}\] \[y''=tan'^{-1}(x)+(x+1)'(1+x^2)^{-1}+(x+1)(1+x^2)'^{-1}\] \[y''=(1+x^2)^{-1}+(1+x^2)^{-1}+(-2x)(x+1)(1+x^2)'^{-2}\]

OpenStudy (amistre64):

quotient rule is fine; i like to use product on negative exponents to make life a bit easier to compute

OpenStudy (anonymous):

i see

OpenStudy (amistre64):

to clean it up we: \[y''=2(1+x^2)^{-1}+(-2x^2-2x)(1+x^2)^{-2}\] \[y''=(2(1+x^2)+(-2x^2-2x))(1+x^2)^{-2}\] \[y''=(2\ \cancel{+2x^2-2x^2}^{\ 0}-2x)(1+x^2)^{-2}\] \[y''=(2-2x)(1+x^2)^{-2}\] y'' = 0 when x=1; and is not undefined for any values so our only option is at x=1 to "test out"

OpenStudy (amistre64):

0 : 0s and undefs <---------------------> - 1 + : curve directions

OpenStudy (amistre64):

well, I got me + and - on the wrong sides, but still; its the point of inflection :)

OpenStudy (anonymous):

so the answer is (-1, 0)?

OpenStudy (anonymous):

nooo i meant (0,1)???

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=2nd+derivative+%28x%2B1%29arctanx i dropped a negative someplace but still the same basic results so inflection is x=1 (x=1,y=) .... need to plug in x=1 into the y= f(x) part to determine its value unless we only have 1 option

OpenStudy (anonymous):

i have the option there, so it's either d or e?

OpenStudy (amistre64):

correct, so when x=1 y=2 arctan(2)

OpenStudy (anonymous):

so (1, pi/2)?

OpenStudy (amistre64):



OpenStudy (amistre64):

i got no idea what angle that is :) wolf to the rescue

OpenStudy (amistre64):

arctan(1) right ....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

k

OpenStudy (amistre64):

2*45^ = 2*pi/4 = pi/2 yep

OpenStudy (anonymous):

thankss again :)